I would like for u to help me check my answers this problem: write 0.7% as a completely reduced fraction.
Found 2 solutions by stanbon, nerdybill:
Answer by stanbon(75887)

write 0.7% as a completely reduced fraction.
---------------------------
% means "per hundred".
--------------------
0.7% = 0.7/100 = 0.007
=============================
Cheers,
Stan H.

Answer by nerdybill(7384)

0.7%
= .7/100
= 7/1000